find gradient of scalar fields

W = 2p (z^2 + 1) cos Φ
1
Expert's answer
2020-04-30T10:41:29-0400
"\nabla W=\nabla (2\rho (z^2 + 1) \cos \phi)"

"\nabla W=2 (z^2 + 1) \cos \phi e_{\rho}+4\rho z\cos \phi e_{z}\\-\left(\rho^{-1}\right)2\rho (z^2 + 1) \sin \phi e_{\phi}"

"\nabla W=2 (z^2 + 1) \cos \phi e_{\rho}+4\rho z\cos \phi e_{z}\\-2 (z^2 + 1) \sin \phi e_{\phi}"